OpenStudy (anonymous):

Prove the identity: (1-sin(x))/(1+sin(x)) = tan(x) + 1/cos(x)

myininaya (myininaya):

\[\frac{1-\sin(x)}{1+\sin(x)} \cdot \frac{1-\sin(x)}{1-\sin(x)}=\frac{1-2\sin(x)+\sin^2(x)}{1-\sin^2(x)}=\frac{1-2 \sin(x)+\sin^2(x)}{\cos^2(x)}\] \[\frac{1}{\cos^2(x)}-2 \frac{\sin(x)}{\cos^2(x)} +\tan^2(x)=\sec^2(x)-2 \frac{\sin(x)}{\cos(x)} \cdot \frac{1}{\cos(x)}+\tan^2(x)\] \[=\sec^2(x)-2 \tan(x) \sec(x)+\tan^2(x)\]

myininaya (myininaya):

i don't see how it can be written as tan(x) + 1/cos(x)

OpenStudy (anonymous):

So are you saying it's not an identity? :|

myininaya (myininaya):

wait i'm looking it is tan(x)+1/cos(x) right?

OpenStudy (anonymous):

tan(x) + (1/cos(x))

myininaya (myininaya):

\[=(\sec(x)-\tan(x))^2\]

myininaya (myininaya):

\[=(1)^2=1\]

OpenStudy (anonymous):

sumpin wrong here

OpenStudy (anonymous):

I've checked wolframalpha and it is indeed an identity that CAN be proven; I'm really stumped though.

myininaya (myininaya):

did i make a mistake?

OpenStudy (anonymous):

here is my proof let \[x=\frac{\pi}{4}\] then \[\frac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}\]is the left hand side, whereas the right hand side is \[1+\sqrt{2}\]

OpenStudy (anonymous):

:|

OpenStudy (anonymous):

and this would be a miracle, since the left hand side is smaller than one, and the right hand side is larger

OpenStudy (anonymous):

so you ought to be stumped, because i cannot be correct.

OpenStudy (anonymous):

sigh..Alright haha

OpenStudy (anonymous):

Thanks anyways

OpenStudy (anonymous):

hey don't believe everything you read in a text book...

OpenStudy (anonymous):

I want to believe this is possible because Wolframlpham says it's a valid identity; but I guess we cannot solve it >_<

OpenStudy (anonymous):

actually wolfram does not say it is an identity. it solved the equation for me here http://www.wolframalpha.com/input/?i=%281-sin%28x%29%29%2F%281%2Bsin%28x%29%29%3Dtan%28x%29+%2B+1%2Fcos%28x%29%2C+

OpenStudy (anonymous):

put "verify" infront of the entire identity

OpenStudy (anonymous):

hold the phone. it is not an identity for sure. i proved it, but wolfram solves the equation for x, which means it is certainly not true for all x

OpenStudy (anonymous):

when i put "verify" it just timed out

OpenStudy (anonymous):

Hmm..So I should conclude that this is not an identity after all?

myininaya (myininaya):

hey why is that one theta bolder type than the other thetas

myininaya (myininaya):

yes we concluded that way up there

OpenStudy (anonymous):

identity means "true for any value of x" like \[2x+x=3x\] is an identity. this one is not

OpenStudy (anonymous):

What if it was a theata instead of "x"

myininaya (myininaya):

\[2 \theta+\theta=3 \theta? \] this is still an identity

myininaya (myininaya):

2p+p=3p is still an identity

OpenStudy (anonymous):

Alright, thanks for the evidence that indeed this is not an identity :)